forces f1 (east) and west f2 are simultaneously applied to 3.0 kg mass. when f2 is east, a= 5.0 m/s east and when f2 is west a=1.0 m/s east. determine the magnitude of f1 and f2


1
Expert's answer
2021-06-29T02:15:03-0400
"F _\n2\n\u200b\n +F _\n1\n\u200b\n =ma _\n1\n\u200b\n =3(5)=15"

"F \n_1\n\u200b\n \u2212F _\n2\n\u200b\n =ma _\n2\n\u200b\n =3(1)=3"

"F _\n1\n\u200b\n =0.5(15+3)=9\\ N"

"F _\n2\n\u200b\n =0.5(15\u22123)=6\\ N"